OpenStudy (anonymous):

What is the gradient of the curve y= 3e^x - 3 where it crosses the y-axis? Find the equation of the tangents to the curve at this point.

OpenStudy (amistre64):

gradient has a few different meanings; but most likely you want the slope of the line in the xy plane for the equation right?

OpenStudy (amistre64):

2 things we need to determine this the derivative of the equation with respect to x; and the value of the equation when x=0; (0,f(0))

OpenStudy (anonymous):

so first thing we need to find dy/dx dy/dx=3e^x at y axis x =0 3*e^0 gradient =3

OpenStudy (precal):

Americans say "slope", British say "gradient"

OpenStudy (anonymous):

american say rise/run :P

OpenStudy (anonymous):

americans*

OpenStudy (anonymous):

haha! thanks and what about the second question?

OpenStudy (anonymous):

Find the equation of the tangents to the curve at this point.

OpenStudy (anonymous):

now plug x=0 in the equation and you will get y=0 gradient of tangent is also 3 applying formula y-0=3(x-0) y=3x

OpenStudy (anonymous):

good night fellows Tc. going to study Chemistry now and my CIES are starting from 8th May wish me luck. take care!